7. Water flows through a pipe of diameter 1.08m at a velocity of 2m/s. If someone puts a nozzle on the end of the pipe, reducing the diameter to 0.35m, at what speed will the water exit the pipe?.

10. A pipe with a radius of 50 cm has water flowing at 5 m/s. If water goes into a new pipe with radius 33 cm, how much will the pressure decrease?

11. A large dam has a hole that is 56 meters from the top. If the water in the reservoir is depending at a rate of 0.33 m/s, how fast is the water exiting the hole?

12. A hand pumped water gun is held at height of 1.3 m above the ground. The water lands 8.4 meters away. What is the exit velocity of the stream? What is the gauge pressure of the gun's reservoir, assuming the velocity of the water is zero and doesn't change height before leaving the reservoir?
Answer #1

7) Using the continuity equation we have

A1v1 = A2v2

pi/4*1.08^2 * 2 = pi/4*0.35^2*v2

v2 = 19.04 m/s

exit speed is 19.04 m/s

10) A1v1 = A2v2

pi*0.5^2*5^2 = pi*0.33^2*v2

v2 = 11.47 m/s

Using the Bernoulli's equation we have

P1 + 0.5pv1^2 = P2 + 0.5pv2^2

P1-P2 = 0.5*p*(v2^2 - v1^2) = 0.5*1000*(11.47^2 - 5^2) = 53377.1 N/m^2